# Robust Equilibria in Shared Resource Allocation via Strengthening Border’s Theorem

### v1

@@ Appendix A Better Equilibrium Guarantees
    Instead of always choosing the allocation probabilities $p_{i}^{S}$ in a very specific way as in Theorem 4.2, let the $p_{i}^{S}$’s be parameters.
We formally give this parameterized mechanism in Mechanism 2.
Input: Per-round budgets $\beta_{i}\in[0,1]$ and slack parameters $\delta_{i}^{T}$
Input: Allocation probabilities $(p_{i}^{S})_{\begin{subarray}{c}i\in[n]\\ S\subseteq[n]\end{subarray}}$ where $(p_{i}^{S})_{i\in[n]}$ is a probability distribution over agents $i\in S$
Endow each agent with $B_{i}[1]=\beta_{i}(1+\delta_{i}^{T})T$ bid tokens;
@@
      Agents submit bids $b_{i}^{t}\in\{0,1\}$;
      Budgets are enforced: $b_{i}^{t}\leftarrow 0$ for each $i$ such that $B_{i}[t]\leq 0$;
      Let $S[t]=\{i:b_{i}^{t}=1\}$ be the set of bidding agents;
      A winner $i^{t}$ is randomly selected from $S[t]$ according to $(p_{i}^{S[t]})_{i\in[n]}$;
      Budgets get updated: $B_{i}[t+1]=B_{i}[t]-b_{i}^{t}$ for every agent $i$;
@@
MECHANISM 2 Generalized Budgeted Border
Regardless of choice of $\beta_{i}$’s and $p_{i}^{S}$’s, there is a similar equilibrium as in Proposition 4.1 and similar equilibrium utility guarantee as in Theorem 4.3. Before stating this formally, let us define some useful terminology. First, we generalize the notion of ideal utility. Adopting terminology from [fikioris2023online], the $\beta$-ideal utility of an agent $i$ is the maximum expected utility they can obtain from a single round if they can obtain the item simply by requesting it, but they are only allowed to request it with probability at most $\beta$. Formally, the (per-round) $\beta$-ideal utility is the following.
Definition A.1 ($\beta$-ideal utility, $\beta$-ideal utility probability function).
@@
Note that the ideal utility, as we previously defined it in Definition 2.1, is just the $\alpha_{i}$-ideal utility.
Recall the notion of $\beta$-aggressive strategy from Definition 4.1, where an agent bids when her value is in the top $\beta$-quantile of her value distribution (subject to her budget constraint). We can also rephrase a $\beta$-aggressive strategy as bidding with probability $(\rho_{i}^{\beta})^{\star}(V_{i}[t])$ (subject to the budget constraint).
For the same reason as in BRB, each agent $i$ playing a $\beta_{i}$-aggressive strategy is an approximate Nash equilibrium. We formally prove the following in Appendix D.
Proposition A.1.
@@
To discuss the utility guarantee at this approximate equilibrium, we define interim allocation probabilities similar to Definition 4.2, except that the interim allocation probabilities $p_{i}$ depend on the per-round budgets $\beta_{i}$. The previous definition of interim allocation probability is a special case where $\beta_{i}=\alpha_{i}$.
Definition A.2.
@@
We now obtain a similar equilibrium utility guarantee as in Theorem 4.3 that we prove formally in Appendix D.
Theorem A.2.
@@
As before, Border’s Theorem gives a clean characterization of which numbers $p_{i}\in[0,1]$ can actually be induced by some allocation probabilities $p_{i}^{S}$. Specifically, the below theorem follows from Border’s Theorem and is analogous to Theorem 5.2, which we explain in Appendix C.
Theorem A.3.
@@
$\sum_{i\in[n]}p_{i}\beta_{i}=1-\prod_{i\in[n]}(1-\beta_{i})$ (8)
@@
$\sum_{i\in I}p_{i}\beta_{i}\leq 1-\prod_{i\in I}(1-\beta_{i}).$ (9)
In general, the principal choose $\beta_{i}$’s and $p_{i}$’s in any way they like that satisfies Theorem A.3. Let’s give a particular suggestion. Suppose we set the $\beta_{i}$’s to be proportional to the fair shares $\alpha_{i}$. We can obtain the following using Theorem A.3, which has a very similar proof to that of Theorem 4.2 that we formally give in Appendix D.
Corollary A.4.
@@
Using the above $p_{i}^{S}$’s in our mechanism, at the equilibrium, each agent $i$ gets expected per-round utility at least $\frac{1-e^{-\gamma}}{\gamma}v^{\star}(\gamma\alpha_{i})$. The function $\frac{1-e^{-\gamma}}{\gamma}$ is decreasing in $\gamma$, while $v^{\star}(\gamma\alpha_{i})$ is nondecreasing in $\gamma$, so if the principal wants to use the above corollary, they could set $\gamma$ appropriately based on the agents’ value distributions to maximize the $\frac{1-e^{-\gamma}}{\gamma}v^{\star}(\gamma\alpha_{i})$. Setting $\gamma=1$ recovers the old mechanism BRB. However, other choices of $\gamma$ may be better if the agents had specific value distributions. For example, the below corollary proved in Appendix D shows that in the symmetric agent case with $\mathrm{Uniform}([0,1])$ value distributions, we can set $\gamma$ such that each agent gets almost all of their $\alpha_{i}$-ideal utility, a result similar to the equilibrium guarantee in [chido].
Corollary A.5.
@@
Appendix B Any Time Guarantees and Exact Nash Equilibrium in an Infinite Time Horizon
B.1 Any Time Guarantees
We can obtain any time utility guarantees by enforcing the budget constraint at any time, i.e., agent $i$ can only bid at most $\alpha_{i}(1+\delta_{i}^{t})t$ times by round $t$ instead of just at end time $T$. We formally describe this mechanism in Mechanism 3 (based off of Generalized Budgeted Border in Appendix A with general per-round budgets $\beta_{i}$).
Input: Per-round budgets $\beta_{i}\in[0,1]$ and slack parameters $\delta_{i}^{t}$
Input: Allocation probabilities $(p_{i}^{S})_{\begin{subarray}{c}i\in[n]\\ S\subseteq[n]\end{subarray}}$ where $(p_{i}^{S})_{i\in[n]}$ is a probability distribution over agents $i\in S$
@@ B.1 Any Time Guarantees
      Agents submit bids $b_{i}^{t}\in\{0,1\}$;
      Budgets are enforced: $b_{i}^{t}\leftarrow 0$ for each $i$ such that $\sum_{s=1}^{t-1}b_{i}^{s}\geq\beta_{i}(1+\delta_{i}^{t})t$;
      Let $S[t]=\{i:b_{i}^{t}=1\}$ be the set of bidding agents;
      A winner $i^{t}$ is randomly selected from $S[t]$ according to $(p_{i}^{S[t]})_{i\in[n]}$;
      Budgets get updated: $B_{i}[t+1]=B_{i}[t]-b_{i}^{t}$ for every agent $i$;
@@
MECHANISM 3 Any Time Budgeted Border
Formally, we obtain the following theorems, proved in Appendix D.
Theorem B.1.
@@
(Here, $p_{i}$ is agent $i$’s interim allocation probability with per-round budgets $(\beta_{k})$ as defined in Definition A.2.)
Theorem B.2.
By running Any Time Budgeted Border with $\beta_{i}=\alpha_{i}$, $\delta_{i}^{t}=\sqrt{\frac{6\ln t}{\alpha_{i}t}}$, and allocation probabilities as guaranteed by Lemmas 6.2 and 6.3, if agent $i$ plays an $\alpha_{i}$-aggressive strategy, regardless of the behavior of other agents $j\neq i$, with probability $1-O(1/\sqrt{t})$, at each time $t$, agent $i$ will obtain utility
@@
B.2 Exact Nash Equilibrium in an Infinite Time Horizon
We show that by running Any Time Budgeted Border in an infinite time horizon, we can make the approximate Nash equilibrium in Theorem 4.3 into an exact Nash equilibrium. Each agent will maximize
@@ B.2 Exact Nash Equilibrium in an Infinite Time Horizon
We prove the following theorem in Appendix D.
Theorem B.3.
By running Any Time Budgeted Border with $\beta_{i}=\alpha_{i}$, $\delta_{i}^{t}=\sqrt{\frac{6\ln t}{\alpha_{i}t}}$, each player $i$ playing a $\beta_{i}$-aggressive strategy is a Nash equilibrium under which
@@
Appendix C Border’s Theorem
C.1 Obtaining Theorems 5.2 and A.3 via a Reduction from Border’s Theorem
We claimed that Theorem 5.2 and more generally, Theorem A.3, is a special case of Border’s Theorem. We shall detail Border’s Theorem and argue why our problem is a special case. Border’s Theorem deals with the setting of selling a single item to bidders via a direct-revelation mechanism. Suppose there are $n$ bidders, each bidder $i$ with type space $\Theta_{i}$. Take a given direct-revelation mechanism. For each reported type profile $\theta=(\theta_{1},\dots,\theta_{n})$, suppose the allocation rule is that bidder $i$ wins with probability $p_{i}^{\theta}$. Define the interim allocation rule to be the mapping $\pi:\bigsqcup_{i}\Theta_{i}\to[0,1]$ where $\pi(\theta_{i})$ denotes the probability that bidder $i$ will win the item conditioned on reporting $\theta_{i}$ assuming others are bidding truthfully. Now suppose instead of taking a given a direct-revelation mechanism and defining the interim allocation rule, we start with an arbitrary function $\pi:\bigsqcup_{i}\Theta_{i}\to[0,1]$. Say that $\pi$ is a feasible interim allocation rule if $\pi$ can arise as an interim allocation rule from an actual direct-revelation mechanism’s allocation rule. Obviously, some functions $\pi$ are not feasible interim allocation rules. For example, $\pi\equiv 1$ is not feasible if $n\geq 2$ because no mechanism can guarantee everyone a probability $1$ of winning regardless of reported type profile.
@@ Appendix C Border’s Theorem
Theorem C.1 (Border’s Theorem).
@@
[…]
The problem of determining which probabilities $p_{i}\in[0,1]$ are interim allocation probabilities induced by allocation probabilities $p_{i}^{S}$ can be seen as a special case of Border’s Theorem. Specifically, assume on a given round in our mechanism that each agent $i$ bids with probability $\beta_{i}$, and we want to set allocation probabilities $p_{i}^{S}$ such that each agent’s probability of winning the item is $p_{i}$ conditioned on bidding. We can think of this problem as a special case of determining whether an interim allocation rule $\pi$ is feasible as follows. Think of each agent $i$ as a bidder with $2$ types: “bidding” and “not bidding” where the probability they have the “bidding” type is $\beta_{i}$. Set $\pi$ to be the function that is 00 on the “not bidding” type and $p_{i}$ on the “bidding” type for agent $i$. Determining whether it is feasible to guarantee each agent probability $p_{i}$ of winning conditioned on bidding is then equivalent to $\pi$ being a feasible interim allocation rule. Using Border’s Theorem to characterize whether $\pi$ is feasible, combined with the additional requirements that we never allocate the item to an agent who is not bidding and we always allocate the item to some agent if at least one agent bids, we obtain the following theorem (a generalization of Theorem 5.2 and a restatement of Theorem A.3).
Theorem C.2.
@@
$\sum_{i\in[n]}p_{i}\beta_{i}=1-\prod_{i\in[n]}(1-\beta_{i})$ (10)
@@
$\sum_{i\in I}p_{i}\beta_{i}\leq 1-\prod_{i\in I}(1-\beta_{i}).$ (11)
We can see that the “only if” direction is obvious as follows. The left-hand side of (10) is the probability that some agent wins the item. The right-hand side of (10) is the probability that some agent bids. These must be equal since we allocate the item whenever at least one agent bids. Now consider (11). The left-hand side is the probability that an agent $i\in I$ wins the item. The right-hand side is the probability that $S$ contains an agent $i\in I$. Since an agent can only win the item if she bids, the left-hand side can be at most the right-hand side.
@@
Proof of Theorem A.3.
Assume (10) and (11) hold. We do a reduction to the problem of determining feasibility of an interim allocation rule. Define a type space $\Theta_{i}=\{0_{i},1_{i}\}$ for each agent $i$ where $\Pr_{\theta_{i}\sim\Theta_{i}}(\theta_{i}=1_{i})=\beta_{i}$. Define $\pi$ by
@@
[…]
by (11). By Theorem C.1, $\pi$ is a feasible interim allocation rule. Find a direct-revelation mechanism with interim allocation rule $\pi$ where $p_{i}^{\theta}$ is the probability that bidder $i$ wins if the reported type profile is $\theta$. Let $p_{i}^{S}=p_{i}^{\theta}$ where $\theta$ has $S=\{i\in[n]:\theta_{i}=1\}$. Notice that $0=\pi(0_{i})$ means that we never allocate the item to an bidder with type $0_{i}$, which implies that $p_{i}^{S}=0$ for $i\notin S$. For $i\in S$, $p_{i}=\pi(1_{i})$ is the probability that we allocate the item to a bidder with type $1_{i}$ conditioned on them having type $1_{i}$. By construction, this is the same conditional probability that we allocate the item to an agent $i$ conditioned on them bidding in Generalized Budgeted Border with allocation probabilities $p_{i}^{S}$. The sum $\sum_{i\in[n]}p_{i}\beta_{i}$ is the probability that bidder $i$ wins the item, and $1-\prod_{i\in[n]}(1-\beta_{i})$ is the probability that some agent $i$ has type $1_{i}$. By (10), the direct-revelation mechanism must allocate the item with probability $1$ to some agent if there is an agent with reported type $1_{i}$. Since $p_{i}^{S}=0$ unless $i\in S$, (10) implies that $\sum_{i\in S}p_{i}^{S}=1$, so $(p_{i}^{S})_{i\in S}$ is indeed a probability distribution over $i\in S$. ∎
C.2 A DMMF Proof of Theorem C.2
Although Theorem C.2 is a special case of Border’s Theorem, to aid intuition, we give novel proof of Theorem C.2 for our special case of allocation to bidding agents by simulating a different request-based allocation mechanism for the same setting.
@@
$\frac{1-\prod_{i\in I}(1-\beta_{i})}{\sum_{i\in I}\alpha_{i}}>1-\prod_{j=1}^{n}(1-\beta_{j}).$ (12)
@@
$\frac{W_{i}[T]}{T}\overset{\mathrm{a.s.}}{\to}\alpha_{i}\left(1-\prod_{j=1}^{n}(1-\beta_{j})\right).$ (13)
Now we give our DMMF proof of Border’s Theorem, which will roughly go as follows. Given interim allocation probabilities $p_{i}$ satisfying the conditions of Theorem A.3, we shall define fair shares $\alpha_{i}$ such that when we run DMMF, the long-run fraction of items that agent $i$ wins is $\beta_{i}p_{i}$. Let $\mu$ be the distribution over $S\subseteq[n]$ where each $i\in S$ independently with probability $\beta_{i}$. We shall show that these fraction items come from the expectation of $\beta_{i}p_{i}^{S}$ over sets $S\sim\mu$ of bidding agents, where $p_{i}^{S}$ is the long-run fraction of times $t$ that agent $i$ would have won had they requested and $S$ were the set of agents who requested at time $t$. These allocation probabilities $p_{i}^{S}$’s will therefore induce the interim allocation probabilities $p_{i}$.
DMMF proof of Theorem C.2.
Necessity of Border’s Criterion for the $p_{i}^{S}$’s to exist was previously argued to be obvious in the remarks after Theorem C.2. For sufficiency, let $(\beta_{i})$ be bid rates and $(p_{i})$ be interim allocation probabilities satisfying Border’s Criterion. We shall assume that (11) holds with strict inequality for $I\subsetneq[n]$; a routine topological argument given in Appendix D extends our proof to the general case.
@@
to be used in DMMF. By (10), the fair shares indeed satisfy $\sum_{i=1}^{n}\alpha_{i}=1$. With these fair shares, by rearrangement, (11) holding with strict inequality and (12) are equivalent. Therefore, the irreducible Markov chain $(Y[t])$ in Theorem C.3 is positive recurrent.
@@
Observe that $\mathcal{S}_{i}[t]$ can be written as a function of $Y[t-1]$. By the pointwise ergodic theorem, for any $S\subseteq[n]$,
@@
Let $S[t]$ be the actual set of agents that request at time $t$. By (13) and the fact that agent $i$ actually wins the item at time $t$ if and only if $S[t]\in\mathcal{S}_{i}[t]$,
@@
$\frac{1}{T}\sum_{t=1}^{T}\operatorname*{\mathbb{E}}_{S\sim\mu}[\boldsymbol{1}\{S\in\mathcal{S}_{i}[t]\}]=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\boldsymbol{1}\{S[t]\in\mathcal{S}_{i}[t]\}].$ (14)

### v2

@@ B Better Equilibrium Guarantees
    Instead of always choosing the allocation probabilities $p_{i}^{S}$ in a very specific way as in Theorem˜3.5, let the $p_{i}^{S}$’s be parameters.
We formally give this parameterized mechanism in Mechanism˜1.
Mechanism 1 Generalized Budgeted Border
0: Per-round budgets $\beta_{i}\in[0,1]$ and slack parameters $\delta_{i}^{T}$.
0: Allocation probabilities $(p_{i}^{S})_{\begin{subarray}{c}i\in[n]\\ S\subseteq[n]\end{subarray}}$ where $(p_{i}^{S})_{i\in[n]}$ is a probability distribution over agents $i\in S$.
 Endow each agent with $B_{i}[1]=\beta_{i}(1+\delta_{i}^{T})T$ bid tokens.
@@
  Agents submit bids $b_{i}^{t}\in\{0,1\}$.
  Budgets are enforced: $b_{i}^{t}\leftarrow 0$ for each $i$ such that $B_{i}[t]\leq 0$.
  Let $S[t]=\{i:b_{i}^{t}=1\}$ be the set of bidding agents.
  A winner $i^{t}$ is randomly selected from $S[t]$ according to $(p_{i}^{S[t]})_{i\in[n]}$.
  Budgets get updated: $B_{i}[t+1]=B_{i}[t]-b_{i}^{t}$ for every agent $i$.
@@
Regardless of choice of $\beta_{i}$’s and $p_{i}^{S}$’s, there is a similar equilibrium as in Proposition˜3.3 and similar equilibrium utility guarantee as in Theorem˜3.6. Before stating this formally, let us define some useful terminology. First, we generalize the notion of ideal utility. Adopting terminology from [fikioris2023online], the $\beta$-ideal utility of an agent $i$ is the maximum expected utility they can obtain from a single round if they can obtain the item simply by requesting it, but they are only allowed to request it with probability at most $\beta$. Formally, the (per-round) $\beta$-ideal utility is the following.
Definition B.1 ($\beta$-ideal utility, $\beta$-ideal utility probability function).
@@
Note that the ideal utility, as we previously defined it in Definition˜2.1, is just the $\alpha_{i}$-ideal utility.
Recall the notion of $\beta$-aggressive strategy from Definition˜3.1, where an agent bids when her value is in the top $\beta$-quantile of her value distribution (subject to her budget constraint). We can also rephrase a $\beta$-aggressive strategy as bidding with probability $(\rho_{i}^{\beta})^{\star}(V_{i}[t])$ (subject to the budget constraint).
For the same reason as in BRB, each agent $i$ playing a $\beta_{i}$-aggressive strategy is an approximate Nash equilibrium. We formally prove the following in subsection˜E.3.
Proposition B.2.
@@
To discuss the utility guarantee at this approximate equilibrium, we define interim allocation probabilities similar to Definition˜3.4, except that the interim allocation probabilities $p_{i}$ depend on the per-round budgets $\beta_{i}$. The previous definition of interim allocation probability is a special case where $\beta_{i}=\alpha_{i}$.
Definition B.3.
@@
We now obtain a similar equilibrium utility guarantee as in Theorem˜3.6 that we prove formally in subsection˜E.3.
Theorem B.4.
@@
As before, Border’s Theorem gives a clean characterization of which numbers $p_{i}\in[0,1]$ can actually be induced by some allocation probabilities $p_{i}^{S}$. Specifically, the below theorem follows from Border’s Theorem and is analogous to Theorem˜4.2, which we explain in section˜D.
Theorem B.5.
@@
(B.1) $\sum_{i\in[n]}p_{i}\beta_{i}=1-\prod_{i\in[n]}(1-\beta_{i})$
@@
(B.2) $\sum_{i\in I}p_{i}\beta_{i}\leq 1-\prod_{i\in I}(1-\beta_{i}).$
In general, the principal choose $\beta_{i}$’s and $p_{i}$’s in any way they like that satisfies Theorem˜B.5. Let’s give a particular suggestion. Suppose we set the $\beta_{i}$’s to be proportional to the fair shares $\alpha_{i}$. We can obtain the following using Theorem˜B.5, which has a very similar proof to that of Theorem˜3.5 that we formally give in subsection˜E.2.
Corollary B.6.
@@
Using the above $p_{i}^{S}$’s in our mechanism, at the equilibrium, each agent $i$ gets expected per-round utility at least $\frac{1-e^{-\gamma}}{\gamma}v^{\star}(\gamma\alpha_{i})$. The function $\frac{1-e^{-\gamma}}{\gamma}$ is decreasing in $\gamma$, while $v^{\star}(\gamma\alpha_{i})$ is nondecreasing in $\gamma$, so if the principal wants to use the above corollary, they could set $\gamma$ appropriately based on the agents’ value distributions to maximize the $\frac{1-e^{-\gamma}}{\gamma}v^{\star}(\gamma\alpha_{i})$. Setting $\gamma=1$ recovers the old mechanism BRB. However, other choices of $\gamma$ may be better if the agents had specific value distributions. For example, the below corollary proved in subsection˜E.2 shows that in the symmetric agent case with $\mathrm{Uniform}([0,1])$ value distributions, we can set $\gamma$ such that each agent gets almost all of their $\alpha_{i}$-ideal utility, a result similar to the equilibrium guarantee in [chido].
Corollary B.7.
@@
C Any Time Guarantees and Exact Nash Equilibrium in an Infinite Time Horizon
C.1 Any Time Guarantees
We can obtain any time utility guarantees by enforcing the budget constraint at any time, i.e., agent $i$ can only bid at most $\alpha_{i}(1+\delta_{i}^{t})t$ times by round $t$ instead of just at end time $T$. We formally describe this mechanism in Mechanism˜1 (based off of Generalized Budgeted Border in section˜B with general per-round budgets $\beta_{i}$).
Mechanism 1 Any Time Budgeted Border
0: Per-round budgets $\beta_{i}\in[0,1]$ and slack parameters $\delta_{i}^{t}$.
0: Allocation probabilities $(p_{i}^{S})_{\begin{subarray}{c}i\in[n]\\ S\subseteq[n]\end{subarray}}$ where $(p_{i}^{S})_{i\in[n]}$ is a probability distribution over agents $i\in S$.
@@ C.1 Any Time Guarantees
  Agents submit bids $b_{i}^{t}\in\{0,1\}$.
  Budgets are enforced: $b_{i}^{t}\leftarrow 0$ for each $i$ such that $\sum_{s=1}^{t-1}b_{i}^{s}\geq\beta_{i}(1+\delta_{i}^{t})t$.
  Let $S[t]=\{i:b_{i}^{t}=1\}$ be the set of bidding agents.
  A winner $i^{t}$ is randomly selected from $S[t]$ according to $(p_{i}^{S[t]})_{i\in[n]}$.
  Budgets get updated: $B_{i}[t+1]=B_{i}[t]-b_{i}^{t}$ for every agent $i$.
@@
Formally, we obtain the following theorems, proved in subsection˜E.4.
Theorem C.1.
@@
(Here, $p_{i}$ is agent $i$’s interim allocation probability with per-round budgets $(\beta_{k})$ as defined in Definition˜B.3.)
Theorem C.2.
By running Any Time Budgeted Border with $\beta_{i}=\alpha_{i}$, $\delta_{i}^{t}=\sqrt{\frac{6\ln t}{\alpha_{i}t}}$, and allocation probabilities as guaranteed by Lemmas˜5.2 and 5.3, if agent $i$ plays an $\alpha_{i}$-aggressive strategy, regardless of the behavior of other agents $j\neq i$, with probability $1-O(1/\sqrt{t})$, at each time $t$, agent $i$ will obtain utility
@@
C.2 Exact Nash Equilibrium in an Infinite Time Horizon
We show that by running Any Time Budgeted Border in an infinite time horizon, we can make the approximate Nash equilibrium in Theorem˜3.6 into an exact Nash equilibrium. Each agent will maximize
@@ C.2 Exact Nash Equilibrium in an Infinite Time Horizon
We prove the following theorem in subsection˜E.4.
Theorem C.3.
By running Any Time Budgeted Border with $\delta_{i}^{t}=\sqrt{\frac{6\ln t}{\alpha_{i}t}}$, each player $i$ playing a $\beta_{i}$-aggressive strategy is a Nash equilibrium under which
@@
D Border’s Theorem
D.1 Obtaining Theorems˜4.2 and B.5 via a Reduction from Border’s Theorem
We claimed that Theorem˜4.2 and more generally, Theorem˜B.5, is a special case of Border’s Theorem. We shall detail Border’s Theorem and argue why our problem is a special case. Border’s Theorem deals with the setting of selling a single item to bidders via a direct-revelation mechanism. Suppose there are $n$ bidders, each bidder $i$ with type space $\Theta_{i}$. Take a given direct-revelation mechanism. For each reported type profile $\theta=(\theta_{1},\dots,\theta_{n})$, suppose the allocation rule is that bidder $i$ wins with probability $p_{i}^{\theta}$. Define the interim allocation rule to be the mapping $\pi:\bigsqcup_{i}\Theta_{i}\to[0,1]$ where $\pi(\theta_{i})$ denotes the probability that bidder $i$ will win the item conditioned on reporting $\theta_{i}$ assuming others are bidding truthfully. Now suppose instead of taking a given a direct-revelation mechanism and defining the interim allocation rule, we start with an arbitrary function $\pi:\bigsqcup_{i}\Theta_{i}\to[0,1]$. Say that $\pi$ is a feasible interim allocation rule if $\pi$ can arise as an interim allocation rule from an actual direct-revelation mechanism’s allocation rule. Obviously, some functions $\pi$ are not feasible interim allocation rules. For example, $\pi\equiv 1$ is not feasible if $n\geq 2$ because no mechanism can guarantee everyone a probability $1$ of winning regardless of reported type profile.
@@ D Border’s Theorem
Theorem D.1 (Border’s Theorem).
@@
$\operatorname*{\mathbb{E}}_{\theta\sim\bigtimes_{i}\Theta_{i}}\left[\sum_{i\in[n]}\pi_{i}(\theta_{i})\boldsymbol{1}_{\{\theta\in\mathcal{T}\}}\right]\leq 1-\prod_{i\in[n]}\left(1-\Pr_{\theta_{i}\sim\Theta_{i}}(\theta_{i}\in\mathcal{T}_{i})\right).$
The problem of determining which probabilities $p_{i}\in[0,1]$ are interim allocation probabilities induced by allocation probabilities $p_{i}^{S}$ can be seen as a special case of Border’s Theorem. Specifically, assume on a given round in our mechanism that each agent $i$ bids with probability $\beta_{i}$, and we want to set allocation probabilities $p_{i}^{S}$ such that each agent’s probability of winning the item is $p_{i}$ conditioned on bidding. We can think of this problem as a special case of determining whether an interim allocation rule $\pi$ is feasible as follows. Think of each agent $i$ as a bidder with $2$ types: “bidding” and “not bidding” where the probability they have the “bidding” type is $\beta_{i}$. Set $\pi$ to be the function that is 0 on the “not bidding” type and $p_{i}$ on the “bidding” type for agent $i$. Determining whether it is feasible to guarantee each agent probability $p_{i}$ of winning conditioned on bidding is then equivalent to $\pi$ being a feasible interim allocation rule. Using Border’s Theorem to characterize whether $\pi$ is feasible, combined with the additional requirements that we never allocate the item to an agent who is not bidding and we always allocate the item to some agent if at least one agent bids, we obtain the following theorem (a generalization of Theorem˜4.2 and a restatement of Theorem˜B.5).
Theorem D.2.
@@
(D.1) $\sum_{i\in[n]}p_{i}\beta_{i}=1-\prod_{i\in[n]}(1-\beta_{i})$
@@
(D.2) $\sum_{i\in I}p_{i}\beta_{i}\leq 1-\prod_{i\in I}(1-\beta_{i}).$
We can see that the “only if” direction is obvious as follows. The left-hand side of (D.1) is the probability that some agent wins the item. The right-hand side of (D.1) is the probability that some agent bids. These must be equal since we allocate the item whenever at least one agent bids. Now consider (D.2). The left-hand side is the probability that an agent $i\in I$ wins the item. The right-hand side is the probability that $S$ contains an agent $i\in I$. Since an agent can only win the item if she bids, the left-hand side can be at most the right-hand side.
@@
Proof 5 (Proof of Theorem˜B.5).
Assume (D.1) and (D.2) hold. We do a reduction to the problem of determining feasibility of an interim allocation rule. Define a type space $\Theta_{i}=\{0_{i},1_{i}\}$ for each agent $i$ where $\Pr_{\theta_{i}\sim\Theta_{i}}(\theta_{i}=1_{i})=\beta_{i}$. Define $\pi$ by
@@
$\operatorname*{\mathbb{E}}_{\theta\sim\bigtimes_{i}\Theta_{i}}\left[\sum_{i\in[n]}\pi_{i}(\theta_{i})\boldsymbol{1}_{\{\theta\in\mathcal{T}\}}\right]=\sum_{i\in I}p_{i}\beta_{i}\leq 1-\prod_{i\in I}(1-\beta_{i})\leq 1-\prod_{i\in[n]}\left(1-\Pr_{\theta_{i}\sim\Theta_{i}}(\theta_{i}\in\mathcal{T}_{i})\right)$
by (D.2). By Theorem˜D.1, $\pi$ is a feasible interim allocation rule. Find a direct-revelation mechanism with interim allocation rule $\pi$ where $p_{i}^{\theta}$ is the probability that bidder $i$ wins if the reported type profile is $\theta$. Let $p_{i}^{S}=p_{i}^{\theta}$ where $\theta$ has $S=\{i\in[n]:\theta_{i}=1\}$. Notice that $0=\pi(0_{i})$ means that we never allocate the item to an bidder with type $0_{i}$, which implies that $p_{i}^{S}=0$ for $i\notin S$. For $i\in S$, $p_{i}=\pi(1_{i})$ is the probability that we allocate the item to a bidder with type $1_{i}$ conditioned on them having type $1_{i}$. By construction, this is the same conditional probability that we allocate the item to an agent $i$ conditioned on them bidding in Generalized Budgeted Border with allocation probabilities $p_{i}^{S}$. The sum $\sum_{i\in[n]}p_{i}\beta_{i}$ is the probability that bidder $i$ wins the item, and $1-\prod_{i\in[n]}(1-\beta_{i})$ is the probability that some agent $i$ has type $1_{i}$. By (D.1), the direct-revelation mechanism must allocate the item with probability $1$ to some agent if there is an agent with reported type $1_{i}$. Since $p_{i}^{S}=0$ unless $i\in S$, (D.1) implies that $\sum_{i\in S}p_{i}^{S}=1$, so $(p_{i}^{S})_{i\in S}$ is indeed a probability distribution over $i\in S$.
D.2 A DMMF Proof of Theorem˜D.2
Although Theorem˜D.2 is a special case of Border’s Theorem, to aid intuition, we give novel proof of Theorem˜D.2 for our special case of allocation to bidding agents by simulating a different request-based allocation mechanism for the same setting.
@@
(D.3) $\frac{1-\prod_{i\in I}(1-\beta_{i})}{\sum_{i\in I}\alpha_{i}}>1-\prod_{j=1}^{n}(1-\beta_{j}).$
@@
(D.4) $\frac{W_{i}[T]}{T}\overset{\mathrm{a.s.}}{\to}\alpha_{i}\left(1-\prod_{j=1}^{n}(1-\beta_{j})\right).$
Now we give our DMMF proof of Border’s Theorem, which will roughly go as follows. Given interim allocation probabilities $p_{i}$ satisfying the conditions of Theorem˜B.5, we shall define fair shares $\alpha_{i}$ such that when we run DMMF, the long-run fraction of items that agent $i$ wins is $\beta_{i}p_{i}$. Let $\mu$ be the distribution over $S\subseteq[n]$ where each $i\in S$ independently with probability $\beta_{i}$. We shall show that these fraction items come from the expectation of $\beta_{i}p_{i}^{S}$ over sets $S\sim\mu$ of bidding agents, where $p_{i}^{S}$ is the long-run fraction of times $t$ that agent $i$ would have won had they requested and $S$ were the set of agents who requested at time $t$. These allocation probabilities $p_{i}^{S}$’s will therefore induce the interim allocation probabilities $p_{i}$.
Proof 6 (DMMF proof of Theorem˜D.2).
Necessity of Border’s Criterion for the $p_{i}^{S}$’s to exist was previously argued to be obvious in the remarks after Theorem˜D.2. For sufficiency, let $(\beta_{i})$ be bid rates and $(p_{i})$ be interim allocation probabilities satisfying Border’s Criterion. We shall first assume that (D.2) holds with strict inequality for $I\subsetneq[n]$; we will then extend our proof to the general case with a topological argument.
@@
to be used in DMMF. By (D.1), the fair shares indeed satisfy $\sum_{i=1}^{n}\alpha_{i}=1$. With these fair shares, by rearrangement, (D.2) holding with strict inequality and (D.3) are equivalent. Therefore, the irreducible Markov chain $(Y[t])$ in Theorem˜D.3 is positive recurrent.
@@
Observe that $\mathcal{S}_{i}[t]$ can be written as a function of Y[t1]Y[t-1]. By the pointwise ergodic theorem, for any $S\subseteq[n]$,
@@
Let $S[t]$ be the actual set of agents that request at time $t$. By (D.4) and the fact that agent $i$ actually wins the item at time $t$ if and only if $S[t]\in\mathcal{S}_{i}[t]$,
@@
(D.5) $\frac{1}{T}\sum_{t=1}^{T}\operatorname*{\mathbb{E}}_{S\sim\mu}[\boldsymbol{1}\{S\in\mathcal{S}_{i}[t]\}]=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\boldsymbol{1}\{S[t]\in\mathcal{S}_{i}[t]\}].$